OpenStudy (anonymous):

solve each equation. show work please.its a fraction just dnt know how to type on here 4d+5/7 +7

OpenStudy (anonymous):

this is not an equation... it doesn't equal anything... this is a function

OpenStudy (anonymous):

you need an = in your equation

OpenStudy (anonymous):

all i can do here is simplify

OpenStudy (anonymous):

i met 4d+5/ =7

OpenStudy (anonymous):

this can be reduced to \[d=\frac{11}{7}\]

OpenStudy (anonymous):

still wrong i met 4d+/7 +7

OpenStudy (anonymous):

or \[\frac{4d+5}{7}=\]

OpenStudy (anonymous):

2nd one

OpenStudy (anonymous):

\[\frac{7(4d+5)}{7}=7(7)\] \[28d+35=49\] \[28d=14\] \[d=2\]

OpenStudy (anonymous):

say you have x/99